A raft foundation subjects its supporting soil to a uniform pressure of 300 kPa. The dimensions of the raft are 6.1m by 15.25m. Determine the vertical stress due to the raft at a depth of 4.58m below it at the central points of the long edges.


1
Expert's answer
2021-05-24T18:22:01-0400


Influence factors for the vertical stress beneath the corner of a rectangular foundation (after

Fadum, 1948).


i) at the center of the raft

z= 4.58 m


m= "\frac{B}{z}= \frac{3.05}{4.58}= 0.67"


n="\frac{L}{z} = \frac{7.63}{4.58}" = 1.67 obtained from the graph.


"l_\alpha = 0. 16"


"4l_\alpha = 0.64"


"\sigma _z= 4l_\alpha P = 0.64 \times 300 = 192 kPa"


ii) at the central point of long edges.


"\Delta\sigma_z= \frac{PBL}{(B+Z)(L+Z)}"


"\Delta\sigma_z = \frac{300\times 15.25 \times 6.1}{(6.1 +4.58)(15.25+4.58)} = 131.65"


"\Delta\sigma_z = 132 kPa"